OpenStudy (anonymous):

How many solutions does the system of equations have? How many solutions does the system of equations have 2x + 3y = 6 y=-2/3x-1

OpenStudy (zarkon):

\[y=-2/3x-1\] \[3(y=-2/3x-1)\] \[3y=-2x-3\] \[2x+3y=-3\]

OpenStudy (anonymous):

so its 4